A parallelogram is formed in R3 by the vectors = (3, 2, –3) and = (4, 1, 5).
The point P = (0, 2, 3).
a. Determine the location of the vertices.
b. Determine the vectors representing the diagonals.
c. Determine the length of the diagonals.
1
Expert's answer
2019-11-25T11:45:46-0500

Solution:a) Let PMNK - a parallelogram.

"\overrightarrow {a}=\overrightarrow {PK}"

"\overrightarrow {PK} (x; y-2; z-3)""x=3; \n\ny-2=2; y=4;\n\nz-3=-3; z=0.\n\n\n\n\nK (3;4;0)"

"\overrightarrow {b}= \overrightarrow {PM}"


"\overrightarrow {PM} (x; y-2; z-3)"

"x=4; y-2=1; y=3; z-3=5; z=8. M(4;3;8)"


"\overrightarrow {a} = \overrightarrow {MN}"


"\overrightarrow {MN} (x-4; y-3; z-8)"


"x-4=3; x=7; \ny-3=2; y=5; \nz-8=-3; z=5. N(7;5;5)"

b)


"\overrightarrow {PN} (7-0;5-2;5-3)=\overrightarrow {PN} (7;3;2)"


"\overrightarrow {MK} (3-4;4-3;0-8)=\overrightarrow {MK} (-1;1;-8)"

c)


"\vert \overrightarrow {PN} \vert= \sqrt {49+9+4} = \sqrt {62}"


"\vert \overrightarrow {MK} \vert = \sqrt {1+1+64} = \sqrt {66}"